# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

### v3

@@ 3.3 Superconducting transition
Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity. We have recently asked Trachenko et al. whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get? Fundamental limit here specifically means the limit in terms of FPCs.
Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (13) Trachenko et al. . $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$. This range of $\lambda$ corresponds to $C=0.1-0.2$. Using Eq. (13), this gives $T_{c}^{u}$ as Trachenko et al.
@@
Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature. Consistent with Eq. (17) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides Flores-Livas et al. (2020); Duan et al. (2017); Pickard et al. (2019). This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above. We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
@@ 3.4 First-order phase transitions: solids, liquids, gases
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. Trachenko (2024) for the functional form of pressure-temperature melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds. This cohesive energy is set by FPCs in Eq. (4). Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq. (11)). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs. This gives the following role of FPCs in setting the states of matter.
@@
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life. First, H2O can exist as liquid water at our typical planetary conditions. Second, carbon-based substances can exist as bonded solidlike structures. Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures would not exist as cohesive states but as the gas states only.
We note that classical entropies of solids and gases contain FPCs explicitly. Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ Landau and Lifshitz (1970), $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq. (12) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
@@
The entropy of the classical gas is Landau and Lifshitz (1970)
@@
With $v\propto a_{\rm B}^{3}$ and $m\propto m_{p}$,
@@
Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref. Trachenko (2024) for a quick review and Refs. Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting Trachenko and Brazhkin (2016); Trachenko (2023c). However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations Lifshitz and Pitaevskii (2006). In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024). The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.
@@ 3.5 Metastable states and chemical reactions
In the previous three sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. Brazhkin (2006) for review). When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006). Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al. .
@@ 4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019). A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007). FPCs would then be calculable from this theory.
@@
In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy. We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of fundamental constants act to increase the entropy. This insight is general and can be applied to different levels of structure in the Universe in Figure 1. This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
@@
We recall that the discussion in Section 3 included long-lived metastable states. We saw that at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy. As recently reviewed Brazhkin (2024), metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)). These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars. In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
@@
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, and its operation are unknown, although several proposals were discussed (see, e.g. Refs. Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013) and Ref. Trachenko (2023d) for review).
@@
Cyclic or oscillatory models of the Universe (see, e.g., Refs. Adams and Laughlin (1997); Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013); Trachenko (2017) and Ref. Kragh (2009) for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe in Ref. Trachenko (2017) at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
@@
We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space and entropy or close to it. Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis in stars and so on). The evolutionary mechanism (with the entropy increase as a selection principle) would imply that structures observed in each cycle are related to the magnitude of fluctuations of FPCs. Large fluctuations would destroy the intricate network of constraints on the values of FPCs needed for different structures in the Universe to exist Adams (2019), resulting in the scarcity of structures and smaller phase space. Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles. There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (similarly to biological evolution, this selection applies to FPCs and FPC-enabled structures existing at the time). This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence. This process involves maximising a property (entropy) as a function of FPCs but within many constraints on their values required to keep different structures and their phase space intact. This invites an input from the optimisation theory.
@@
An example related to this discussion and involving constraints is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), breaking many constraints required for key processes of structure formation and disrupting these processes Barrow (2003); Adams (2019). This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs). We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on. Given the scale of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
@@ 5 Discussion and summary
Approaching the challenge to explain fundamental constants started with understanding the role these constants play. This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties. This has enriched condensed matter physics and deepened our understanding of its key phenomena.
@@
This list of effects and phenomena in condensed matter can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. Adams (2019) for a review of these challenges).
If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe. The insight that FPCs act to increase the entropy is one such example. This and similar insights can provide useful links between different areas of physics and open up new discussions.
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and superconductivity Trachenko et al. . Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.

### v4

@@ 3.3 Superconducting transition
Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity. We have recently asked [32] whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get? Fundamental limit here specifically means the limit in terms of FPCs.
Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (13) [32]. $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$. This range of $\lambda$ corresponds to $C=0.1-0.2$. Using Eq. (13), this gives $T_{c}^{u}$ as [32]
@@
Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature. Consistent with Eq. (17) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides [39, 40, 41]. This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above. We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
@@ 3.4 First-order phase transitions: solids, liquids, gases
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. [42] for the functional form of pressure-temperature melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy [43] because microscopically, melting involves breaking chemical bonds. This cohesive energy is set by FPCs in Eq. (4). Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems [43, 44] (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq. (11)). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs. This gives the following role of FPCs in setting the states of matter.
@@
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life. First, H2O can exist as liquid water at our typical planetary conditions. Second, carbon-based molecules and compounds can exist as cohesive (bonded) structures. Much larger cohesive energy due to different values of FPCs would result in H2O existing as a solid only and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would result in H2O and carbon-based structures not existing as cohesive states but as gas states only.
We note that classical entropies of solids and gases contain FPCs explicitly. Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ [36], $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq. (12) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
@@
The entropy of the classical gas is [36]
@@
Using $v\propto a_{\rm B}^{3}$, Eq. (3) and $m\propto m_{p}$ gives
@@
Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref. [42] for a quick review and Refs. [45, 46] for a more detailed discussion). Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting [45, 46]. However, the energy, the primary property in statistical physics [36], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations [47]. In liquids, these excitations are gapped phonons and particle jumps [45, 46, 48, 49, 50, 51]. The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.
@@ 3.5 Metastable states and chemical reactions
In the previous three sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. [52] for review). When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable [52]. Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) [53].
@@ 4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation. Theories and observations of this variation were previously discussed [54, 55, 56, 57, 58, 59, 60, 6, 13, 61]. Anthropic arguments use this variation to rationalise the observed values of FPCs [9, 10, 13]. A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) [33] and guided by a unique design and an exquisite symmetry [62]. FPCs would then be calculable from this theory.
@@
In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases at given environmental parameters, increase the entropy. We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of FPCs give rise to entropy increase.
We discussed the entropy increase due to FPCs in condensed matter systems, however this increase is a general effect and can be applied to different length and energy scales in the Universe on general grounds. This is supported by a more specifc observation. The temperature of transitions and reactions at different energy scales is related to the energy of interaction between constituents and the relevant cohesive energy at each scale. These energies depend on FPCs, implying that FPCs enable transitions and reactions at those scales as they do in condensed matter. Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy. Hence we expect that FPCs act to increase the entropy at different levels of structure in the Universe in Figure 1.
The assertion of the entropy increase includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed [36]. We also note that the environmental parameters can be variable in some processes. For example, nuclear synthesis reactions in a forming star depend on (environmental) pressure and temperature, which, in turn, depend on the mass of collapsing interstellar gas. This variability can affect transitions and reactions and the properties of emergent structures. In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions. Nevertheless, the transitions and reactions are enabled or disabled by FPCs at varying environmental parameters as they are at fixed parameters.
@@
We recall that the discussion in Section 3 included long-lived metastable states. We saw that at given environmental parameters, the collection of these states increase the phase space and entropy. As recently reviewed [63], metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe [63]). These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which are sensitive to the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
@@
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, and its operation are unknown, although several proposals were discussed (see, e.g. Refs. [64, 65, 66, 67, 68] and Ref. [61] for review).
@@
Cyclic or oscillatory models of the Universe (see, e.g., Refs. [69, 64, 65, 66, 67, 68, 70] and Ref. [71] for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe in Ref. [70] at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
@@
We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it. Adams concludes [13] that our Universe is not the most abundant and complex possible: somewhat different FPCs would give more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis and so on). Another interesting observation in Ref. [13] is that our Universe does not lie at the centre of allowed range of FPCs. The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations. We first note that this selection applies to FPCs and FPC-enabled structures existing at the time (as is the case in biological evolution) rather than to currently observed structures. Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist [13], resulting in scarcity of structures and smaller phase space. Smaller fluctuations maintain the continuity of structures in different cycles. As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along this sequence. Third, maximising a property (entropy) as a function of FPCs is subject to many constraints on their values. This invites an input from the optimisation theory.
@@
An example related to this discussion and involving constraints is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs [72], one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), breaking many constraints required for key processes of structure formation and disrupting these processes [6, 13]. This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times [6, 13] (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs). We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe [6, 13], identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity [36] and so on. Given the scale of the problem we face in understanding the origin of fundamental constants [6, 15, 14], making new observations and asking new questions is useful for progress.
@@ 5 Discussion and summary
Approaching the challenge to explain fundamental constants started with understanding the role these constants play. This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures [6, 7, 8, 9, 10, 11, 12, 13, 1]. Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties. This has enriched condensed matter physics and deepened our understanding of its key phenomena.
@@
This list of effects and phenomena in condensed matter can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. [13] for a review of these challenges).
If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe. The insight that FPCs promote entropy increase is one such example. This and similar insights can provide useful links between different areas of physics and open up new discussions.
Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models. These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30] and superconductivity [32]. Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.